2 x + 2 |x| >=2√2 Solve.. What is the answer .Please find and reply me fast

2x + 2|x| >=2√2
Solve..
What is the answer .Please find and reply me fast

Grade:

1 Answers

Arun
5 years ago
Dear student
 
if x >0
2^x + 2^x > = 2 * sqrt(2)
2^x > = sqrt(2)
x > = ½
if x
2^x + 2^ –x > = 2 * sqrt(2)
2^2x + 1 > = 2^(x+1)  * sqrt (2)
assume 2^x = t and then it will be a quadratic equation. now you can solve it further